# Error correction of transversal controlled-NOT gates for scalable surface code computation

### v1

@@ Appendix D Logical computation using lattice surgery
Figure 9: (a) Setup, and (b) spacetime volume of an $XX$ logical measurement conducted using lattice surgery, with (i) the product of stabilizers denoting the $XX$ measurement outcome, and (ii) an error that will be misidentified without the buffer measurement rounds. (c) Spacetime decoding volume for a logical CNOT done with lattice surgery.
@@
We illustrate the process for a joint-Pauli $XX$ measurement done via lattice surgery on two surface codes in Fig. 9(a)-(b). The system is set up as follows: two initially disconnected surface codes have their individual stabilizers measured for $O(d)$ rounds. A bridging region of width $b$ between the $X$-logical edges of these codes is initialized with unentangled qubits in |$\ket{0}$ (Fig. 9(a)). To begin the joint-parity (JP) measurement, surface-code checks are measured over the entire region comprising of the two logical qubits and bridging region. This projects the two surface codes into a joint logical subspace.
@@
In the absence of errors, all check measurements in the logical region, and $Z$ stabilizer measurements in the bridging region return $+1$ outcomes. $X$ stabilizers in the bridging region are initially unfixed and yield random outcomes, but their combined product is the $XX$ logical measurement outcome (Fig. 9(b)(i)). To achieve fault-tolerance to measurement errors that may lead to misidentification of the $XX$ measurement result, JP measurements are done for $O(d)$ rounds. After these JP rounds, the bridging region qubits are measured in the $Z$ basis, and the surface codes are once again measured independently. The spacetime decoding volume of this process is shown in Fig. 9(b).
@@
Extending this strategy, a logical CNOT gate between a control surface code $C$ and a target surface code $T$ can be performed using a similar JP-measurement based circuit. This operation uses a logical $ZZ$ measurement on $C$ and a logical ancilla patch, followed by a logical $XX$ measurement on the ancilla and $T$. In practice, the spacetime volume of this set of operations looks like Fig. 9(c) where $O(2d)$ round of gates along with one logical ancilla are used to perform the complete gate. Note that single-control $n$-target CNOT gates can be carried out simultaneously by using a large ancillary region that can support $X^{\otimes n}$ measurements [fowler_low_2018, litinski_lattice_2018]. This CNOT operation has the same threshold as a SCQM, and the same logical error rate scaled up to known constant prefactors (discussed in App. E). This is verified via numerical simulations in Fig. 4.
Appendix E Error correction in lattice surgery
@@
We expand on decoding and correction of errors during the $XX$ joint parity measurement discussed in App. D. Close observation of the ‘legs’ of the H in Fig. 9(b) – representing pre-JP stabilizer measurement rounds on the individual surface codes, hereon referred to as buffer rounds - show that they are equivalent to the spacetime volume of individual SCQM experiments (up to time boundaries that connect them to the JP rounds). The bridging region in the joint parity rounds is a representation of the stability experiment [gidney_stability_2022].The stability experiment can be interpreted as a space-time rotated version of the SCQM. It tracks a stabilizer product - here, the product of $X$ stabilizers in the bridging region - and succeeds if the predicted value based on measurement outcomes matches the true value on the underlying qubits.
@@
As a combination of variously orientated SCQMs, we can observe that the complete decoding graph of the logical $XX$ measurement is essentially a trivially extended SCQM. As a result, in lattice surgery, we preserve the threshold of an equivalent SCQM. This has previously been demonstrated in Ref. [chamberland_universal_2022] using a matching decoder. Further, the total logical error rate of a lattice surgery operation can be expressed in terms of its logical spacetime surface area (LSSA) in comparison to an SCQM’s LSSA, a concept we discuss in App. F.
@@
An additional note of interest is the choice to use buffer rounds at all; we connect this to the notion of $W$ in the main text. Technically, these legs of the ‘H’ shaped spacetime volume are not part of the JP measurement itself. Indeed, in the ideal scenario of perfect state preparation of the disjoint surface codes, or surface code readout immediately after the JP rounds, they are not required. However, measurement errors in the original logical patches can create defects that can be misidentified as faulty joint-parity measurement errors had measurement data from the buffer rounds not been used. Fig. 9(b)(ii) exemplifies this: if the decoding window does not extend below the first JP round, the lone defect above the window boundary is misidentified as arising from a JP $X$ stabilizer measurement error. Thus, we use $W$ buffer rounds of decoding data for fault-tolerance against these errors. Like the tCNOT, these buffer rounds can be part of other non-commuting JP measurements - inclusion of their measurement data is only necessary for decoding and they do not actually need to be perfectly disjoint rounds on individual surface codes (see Fig. 9(c)).
Appendix F Logical surface areas for SC operations
@@
We now move onto a 2SCQM, i.e. a memory experiment on two surface codes for $2d$ rounds, the logical spacetime surface area (LSSA) of which is:.
@@
$\frac{XX_{2SC+3d}}{I_{2SCQM}}=1.375+\frac{b}{4d}$ (3)
This is verified numerically in Fig. 10 which uses $b=1$.
Figure 10: Ratio of logical failures rates for a $XX$ measurement using lattice surgery with $b=1$ to a 2SCQM. Each point uses $10^{5}$ samples, discarding points with failure rates $\leq 10^{-4}$. Error bars represent the 95% confidence interval. A horizontal line is drawn at the lower bound of $1.375$.
We can similarly calculate the LSSA for a logical CNOT between two surface codes using lattice surgery. As described in App. E, this operation takes a total of $4d$ rounds. Its LSSA ${CX}_{2SC+4d}$ is:
@@
$\frac{CX_{2SC+4d}}{I_{2SCQM}}=2.25+\frac{b}{2d}$ (4)
F.1 LSSAs for tCNOTs
We now calculate the magnification of error rates induced by ordered decoding at low $p$. The overall failure rate of the tCNOT is the probability that the decoder applies a logically incorrect update to any of the control-$X$, control-$Z$, target-$X$, or target-$Z$ decoding graphs. For a 2SCQM, these four rates are equivalent. For a tCNOT with ordered decoding, we can deduce that the dependent graphs have a 50% increase in logical failure rates compared to a 2SCQM because they use their own LSSA combined with half the LSSA of their corresponding independent subgraph. The independent subgraphs remain unaffected. Thus, the logical rate amplifier is: $\frac{2+2\times{1.5}}{2+2}=1.25$
Figure 11: Ratio of logical failures rates for a tCNOT with ordered decoding to a 2SCQM. Each point uses $10^{5}$ samples, discarding points with failure rates $\leq 10^{-4}$. Error bars represent the 95% confidence interval. Horizontal lines are drawn at $1$ and $1.25$.
In Fig. 11, we verify these calculations. We find the ratio $\mathcal{R}$ of the numerically obtained logical failure rates for a tCNOT corrected using ordered decoding to that of a 2SCQM for various system sizes. By virtue of the ambiguous errors, along with the errors introduced within the faulty physical CNOT gates that implement the tCNOT, we observe that $\mathcal{R}$ is consistently higher than 1 and seems to saturate below threshold. By fitting to a saturation function, we find a saturation $\mathcal{R}\approx 1.25$ at low $p$, represented by the dark green horizontal line. This agrees with the analytical prediction.
F.2 Multi-target CNOTs
@@
$\frac{{CX_{\text{LS-n}}}}{I_{(n+1)SC+2d}}=\frac{21+15n+\lceil n/2\rceil 8bd}{8(n+1)}$ (5)
@@
$\frac{{CX_{\text{t-n}}}}{I_{(n+1)SC+2d}}=\frac{(n+1)+n\times 1.5+(1+0.5n)}{2(n+1)}$ (6)
Appendix G An isolated LS-CNOT vs a tCNOT
@@
Lattice surgery $d^{2}+2bd$ $2d$ $O(\left(10d^{3}+2bd^{2}\right)^{3})$ Y $2.25+b/2d$ [App. F]
tCNOT + ordered decoding 0 $d$ $2\times O(\left(2d^{3}\right)^{3})$ N $\sim 1.25$ [App. F]
@@
Table 1: A preliminary overhead comparison for an isolated fault-tolerant logical CNOT performed via lattice surgery vs. transversally with ordered decoding. We use a MWPM decoding subroutine, resulting in both CNOT strategies having a threshold of $1\%$. We benchmark the number of additional qubits and measurement rounds needed for both strategies, along with the decoding time complexity and total logical error rate in comparison to a 2SCQM.
@@
Here we discuss our results benchmarking the performance of a fault-tolerant tCNOT against a lattice-surgery-CNOT, in a setting where state preparation and measurement do not immediately occur before and after the logical operation. We highlight that this comparison is quite artificial, given that (1) in practice for transversal circuits, it may be more advantageous to perform $m>1$ consecutive alternating tCNOTs in a logical decoding block, and (2) in practice for joint-measurement based circuits, Clifford gates can be time-efficiently compiled [fowler_time-optimal_2013]. A more realistic analysis is left to future work.
@@
We find that for a tCNOT, the total logical error rate is only marginally amplified compared to a 2SCQM experiment. However, it more than doubles for a LS-CNOT on account of that fact that it needs two logical Pauli measurements (see Appendix. F for an analysis). Further, the lattice surgery approach requires an additional ancilla surface code patch, as well as bridging regions of uninitialized qubits. This results in a total additional qubit overhead of $d^{2}+2bd$, where $b$ is the width of the bridging region of qubits. $b$ is dependant on the architecture chosen. For movable architectures, this can be $O(1)$, but for fixed architectures this is generally the separation of the surface codes patches. These extra qubits are not needed for a tCNOT. Using a MWPM subroutine, the time overhead for ordered decoding of the tCNOT scales as $2\times O(\left(2d^{3}\right)^{3})$. The factor of two arises since the independent and dependent graphs are decoded serially. This same quantity is $O(\left(10d^{3}+2bd^{2}\right)^{3})$ for lattice surgery, implying that despite the inherent latency of ordered decoding, a correction $\mathcal{C}$ may be identified faster than an equivalent lattice surgery instance. The results of this comparison are summarized in Table  1.
Appendix H The biased erasure noise model
@@
Appendix I Details of numerical simulations
The simulations based on MWPM decoders for the SCQM and transversal CNOT operations were implemented in C++ using the Blossom algorithm [edmonds_paths_1965, kolmogorov_blossom_2009], combined with Djikstra to account for adjusted edge weights arising from erasures. The code for these is available at https://github.com/kaavyas99/CSS-tCNOT-decoders. The corresponding HUF-based simulations were implemented in Stim [gidney_stim_2021], and decoded using MWPF [wu_wuyue16pkugmailcom_mwpf_nodate]. A second implementation of ordered decoding in Stim using PyMatching [gidney_stim_2021, higgott_sparse_2023] was additionally developed for further testing.
In the noise model used for simulations, two-qubit gate errors are uniformly chosen at random from $\{I,X,Y,Z\}^{\otimes 2}/\{I\otimes I\}$ at a rate $p$. This model also applies to the transversal CNOT gates themselves. State preparation, measurement, idling, and single-qubit gate errors are not included [fowler_surface_2012].

### v2

@@ Appendix E Logical computation using lattice surgery
Figure A3: (a) Setup, and (b) spacetime volume of an $XX$ logical measurement conducted using lattice surgery, with (i) the product of stabilizers denoting the $XX$ measurement outcome, and (ii) an error that will be misidentified without the buffer measurement rounds. (c) Spacetime decoding volume for a logical CNOT done with lattice surgery.
@@
We illustrate the process for a joint-Pauli $XX$ measurement done via lattice surgery on two surface codes in Fig. A3(a)-(b). The system is set up as follows: two initially disconnected surface codes have their individual stabilizers measured for $O(d)$ rounds. A bridging region of width $b$ between the $X$-logical edges of these codes is initialized with unentangled qubits in |$\ket{0}$ (Fig. A3(a)). To begin the joint-parity (JP) measurement, surface-code checks are measured over the entire region comprising of the two logical qubits and bridging region. This projects the two surface codes into a joint logical subspace.
@@
In the absence of errors, all check measurements in the logical region, and $Z$ stabilizer measurements in the bridging region return $+1$ outcomes. $X$ stabilizers in the bridging region are initially unfixed and yield random outcomes, but their combined product is the $XX$ logical measurement outcome (Fig. A3(b)(i)). To achieve fault-tolerance to measurement errors that may lead to misidentification of the $XX$ measurement result, JP measurements are done for $O(d)$ rounds. After these JP rounds, the bridging region qubits are measured in the $Z$ basis, and the surface codes are once again measured independently. The spacetime decoding volume of this process is shown in Fig. A3(b).
@@
Extending this strategy, a logical CNOT gate between a control surface code $C$ and a target surface code $T$ can be performed using a similar JP-measurement based circuit. This operation uses a logical $ZZ$ measurement on $C$ and a logical ancilla patch, followed by a logical $XX$ measurement on the ancilla and $T$. In practice, the spacetime volume of this set of operations looks like Fig. A3(c) where $O(2d)$ round of gates along with one logical ancilla are used to perform the complete gate. Note that single-control $n$-target CNOT gates can be carried out simultaneously by using a large ancillary region that can support $X^{\otimes n}$ measurements [fowler_low_2018, litinski_lattice_2018]. This CNOT operation has the same threshold as a SCQM, and the same logical error rate scaled up to known constant prefactors (discussed in App. F). This is verified via numerical simulations in Fig. 4.
Appendix F Error correction in lattice surgery
@@
We expand on decoding and correction of errors during the $XX$ joint parity measurement discussed in App. E. Close observation of the ‘legs’ of the H in Fig. A3(b) – representing pre-JP stabilizer measurement rounds on the individual surface codes, hereon referred to as buffer rounds - show that they are equivalent to the spacetime volume of individual SCQM experiments (up to time boundaries that connect them to the JP rounds). The bridging region in the joint parity rounds is a representation of the stability experiment [gidney_stability_2022].The stability experiment can be interpreted as a space-time rotated version of the SCQM.
@@
As a combination of variously orientated SCQMs, we can observe that the complete decoding graph of the logical $XX$ measurement is essentially a trivially extended SCQM. As a result, in lattice surgery, we preserve the threshold of an equivalent SCQM. This has previously been demonstrated in Ref. [chamberland_universal_2022] using a matching decoder. Further, the total logical error rate of a lattice surgery operation can be expressed in terms of its logical spacetime surface area (LSSA) in comparison to an SCQM’s LSSA, a concept we discuss in App. G.
@@
An additional note of interest is the choice to use buffer rounds at all; we connect this to the notion of $W$ in the main text. Technically, these legs of the ‘H’ shaped spacetime volume are not part of the JP measurement itself. Indeed, in the ideal scenario of perfect state preparation of the disjoint surface codes, or surface code readout immediately after the JP rounds, they are not required. However, measurement errors in the original logical patches can create defects that can be misidentified as faulty joint-parity measurement errors had measurement data from the buffer rounds not been used. Fig. A3(b)(ii) exemplifies this: if the decoding window does not extend below the first JP round, the lone defect above the window boundary is misidentified as arising from a JP $X$ stabilizer measurement error. Thus, we use $W$ buffer rounds of decoding data for fault-tolerance against these errors. Like the tCNOT, these buffer rounds can be part of other non-commuting JP measurements - inclusion of their measurement data is only necessary for decoding and they do not actually need to be perfectly disjoint rounds on individual surface codes (see Fig. A3(c)).
Appendix G Logical surface areas for SC operations
@@
We now describe a 2SCQM, i.e. a memory experiment on two surface codes for $2d$ rounds, the logical spacetime surface area (LSSA) of which is:.
@@
$\frac{XX_{2SC+3d}}{I_{2SCQM}}=1.375+\frac{b}{4d}$ (A2)
This is verified numerically in Fig. A4 which uses $b=1$.
Figure A4: Ratio of logical failures rates for a $XX$ measurement using lattice surgery with $b=1$ to a 2SCQM. Each point uses $10^{5}$ samples, discarding points with failure rates $\leq 10^{-4}$. Error bars represent the 95% confidence interval. A horizontal line is drawn at the lower bound of $1.375$.
We can similarly calculate the LSSA for a logical CNOT between two surface codes using lattice surgery. As described in App. F, this operation takes a total of $4d$ rounds. Its LSSA ${CX}_{2SC+4d}$ is:
@@
$\frac{CX_{2SC+4d}}{I_{2SCQM}}=2.25+\frac{b}{2d}$ (A3)
G.1 LSSAs for tCNOTs
We now calculate the magnification of error rates induced by ordered decoding at low $p$. The overall failure rate of the tCNOT is the probability that the decoder applies a logically incorrect update to any of the four $\{G_{CX},G_{CZ},G_{TX},G_{TX}\}$ decoding subgraphs. For a 2SCQM, these four rates are equivalent. For a tCNOT with ordered decoding, we take the dependent subgraphs to have a 50% increase in logical failure rates compared to a 2SCQM because they use their own LSSA combined with half the LSSA of their corresponding independent subgraph. The independent subgraphs remain unaffected. Thus, the logical rate amplifier is: $\frac{2+2\times{1.5}}{2+2}=1.25$
Figure A5: Ratio of logical failures rates for a tCNOT with ordered decoding to a 2SCQM. Each point uses $10^{5}$ samples, discarding points with failure rates $\leq 10^{-4}$. Error bars represent the 95% confidence interval. Horizontal lines are drawn at $1$ and $1.25$.
In Fig. A5, we verify these calculations. We find the ratio $\mathcal{R}$ of the numerically obtained logical failure rates for a tCNOT corrected using ordered decoding to that of a 2SCQM for various system sizes. By virtue of incorrectly propagated logical updates, ambiguous errors, and errors introduced within the faulty physical CNOT gates that implement the tCNOT, we observe that $\mathcal{R}$ is consistently higher than 1 and seems to saturate below threshold. By fitting to a saturation function, we find a saturation $\mathcal{R}\approx 1.25$ at low $p$, represented by the dark green horizontal line. This agrees with the preliminary analytical prediction above.
G.2 Multi-target CNOTs
@@
$\frac{{CX_{\text{LS-n}}}}{I_{(n+1)SC+2d}}=\frac{21+15n+\lceil n/2\rceil 8bd}{8(n+1)}$ (A4)
@@
$\frac{{CX_{\text{t-n}}}}{I_{(n+1)SC+2d}}=\frac{(n+1)+n\times 1.5+(1+0.5n)}{2(n+1)}$ (A5)
Appendix H An isolated LS-CNOT vs a tCNOT
@@
Lattice surgery $d^{2}+2bd$ $2d$ $O(\left(10d^{3}+2bd^{2}\right)^{3})$ Y $2.25+b/2d$ [App. G]
tCNOT + ordered decoding 0 $d$ $2\times O(\left(2d^{3}\right)^{3})$ N $\sim 1.25$ [App. G]
@@
Table A1: A preliminary overhead comparison for an isolated fault-tolerant logical CNOT performed via lattice surgery vs. transversally with ordered decoding. We use a MWPM decoding subroutine, resulting in both CNOT strategies having a threshold of $1\%$. We benchmark the number of additional qubits and measurement rounds needed for both strategies, along with the decoding time complexity and total logical error rate in comparison to a 2SCQM.
@@
Here we discuss our results benchmarking the performance of a fault-tolerant tCNOT against a lattice-surgery-CNOT, in a setting where state preparation and measurement do not immediately occur before and after the logical operation. We highlight that this comparison is quite artificial, given that (1) in practice for transversal circuits, it may be more advantageous to use $m,g>1$ tCNOTs in a logical decoding block (see Sec. III), and (2) in practice for joint-measurement based circuits, Clifford gates can be time-efficiently compiled [fowler_time-optimal_2013]. A more realistic analysis is left to future work.
@@
We find that for a tCNOT, the total logical error rate is only marginally amplified compared to a 2SCQM experiment. However, it more than doubles for a LS-CNOT on account of that fact that it needs two logical Pauli measurements (see Appendix. G for an analysis). Further, the lattice surgery approach requires an additional ancilla surface code patch, as well as bridging regions of uninitialized qubits. This results in a total additional qubit overhead of $d^{2}+2bd$, where $b$ is the width of the bridging region of qubits. $b$ is dependant on the architecture chosen. For movable architectures, this can be $O(1)$, but for fixed architectures this is generally the separation of the surface codes patches. These extra qubits are not needed for a tCNOT. Using a MWPM subroutine, the time overhead for ordered decoding of the tCNOT scales as $2\times O(\left(2d^{3}\right)^{3})$. The factor of two arises since the independent and dependent graphs are decoded serially. This same quantity is $O(\left(10d^{3}+2bd^{2}\right)^{3})$ for lattice surgery, implying that despite the inherent latency of ordered decoding, a correction $\mathcal{C}$ may be identified faster than an equivalent lattice surgery instance. The results of this comparison are summarized in Table  A1.
Appendix I Details of numerical simulations
The simulations based on MWPM decoders for the SCQM and transversal CNOT operations were implemented in C++ using the Blossom algorithm [edmonds_paths_1965, kolmogorov_blossom_2009], combined with Djikstra to account for adjusted edge weights arising from erasures. The code for these is available at https://github.com/kaavyas99/CSS-tCNOT-decoders. The corresponding HUF-based simulations were implemented in Stim [gidney_stim_2021], and decoded using the MWPF implementation of HUF [wu_wuyue16pkugmailcom_mwpf_nodate]. A second implementation of ordered decoding in Stim using PyMatching [gidney_stim_2021, higgott_sparse_2023] was additionally developed for further testing.
In the noise model used for simulations, two-qubit gate errors are uniformly chosen at random from $\{I,X,Y,Z\}^{\otimes 2}/\{I\otimes I\}$ at a rate $p$. This model also applies to the physical CNOT gates used in the logical tCNOT. State preparation, measurement, idling, and single-qubit gate errors are not included [fowler_surface_2012].
Appendix J The biased erasure noise model